# Your copula is a classifier in disguise: classification-based copula density estimation

### v1

@@ Appendix C Practical details
C.1 High dimensional experiments
The two high-dimensional datasets used are the Digits dataset (Alimoglu and Alpaydin,, 1997) and MNIST (LeCun,, 1998). They both show handwritten digits in $8\cdot 8$ or $28\cdot 28$ format. As both these datasets represent complex dependencies, as measured using Lemma 3.2 from the main text, we use a GG+ratio copula model. In Figure 4, we show random samples from the ratio copula, the simplified vine copula, the IGC and the Gaussian copula as well as true test data. We show samples from the same copula classes fitted on the MNIST data in Figure 5. Samples shown are obtained from the copula model and then transformed back to the data space using the empirical CDF as described in the preprocessing paragraph. We note that the samples shown are random across tiles, meaning that we are not showing how well a model imitates a given target digit. We are exploring the learned distribution implied by the model after training on the data samples, doing so by showing a collection of uncurated random samples from any given model.

### v2

@@ Appendix C PRACTICAL DETAILS
C.1 Additional experiments on losses for tail modelling
In Menon and Ong [2016], authors propose losses that emphasize high density-ratio values. In copulas, under our ratio copula perspective, this corresponds to extreme tail dependence.
We exemplify the effect of different losses which place more or less emphasis on the tails for estimating ratio copulas. The Noise Contrastive Estimation (NCE) loss of Gutmann and Hyvärinen [2012] we use emphasizes training according to a weight in terms of the copula value $c(\mathcal{P}(\mathbf{x}))$ as $w(c)=1/(c^{2}+c)$, while the exponential loss has $w(c)=1/c^{(3/2)}$ and the least squares importance fitting (LSIF) loss [Kanamori et al., 2009] has $w(c)=1$, both latter losses placing a greater focus on tails. We illustrate this behaviour on a two-dimensional example shown in Figure 5. For observed data on the first row, we fit the same ratio copula model using different loss functions, showing the resulting classifiers and log-likelihoods on the left and right columns respectively. The exponential loss better captures tail events (boxed in blue), with sharper densities and better separation of observation ”rays”, while NCE conflates them. The higher LL values given to data-regions highlight the advantage of the exponential loss. The LSIF loss is harder to optimize which makes its density more diffuse. We hereby demonstrate another advantage of connecting copulas to DRE, allowing practitioners to make use of already developed tools, in this instance to improve tail modelling.
Figure 5: Same Ratio copula model trained with losses that emphasise tail behaviours. The exponential loss obtains a better separation of likelihood regions in the tails (boxed in blue) while the NCE loss conflates tail events and the LSIF loss is too diffuse.
C.2 Experiment on bivariate copulas from pictures in Figure 2
The data to fit was obtained by normalising the pixel values of a monochromatic picture, to sum up to 1 across the image. We then sample with replacement following the probabilities given by the brightness of each pixel and add minute Gaussian noise to dequantize samples. This constitutes our dataset on $\mathbb{R}^{2}$, for which we estimate marginals (with empirical CDFs) and use them to finally obtain copula samples (the plots from the first column, with white dots for each sample) on $[0,1]^{2}$ used to fit copula models. Notice how the letters of the bottom row have different shades of white: this is because of the transformation to the copula space. Our model recovers these nuances, see the horizontal bars of each ”T” which have a larger density than the vertical bars. In Figure 2, we display on $[0,1]^{2}$, going from left to right: the observed dataset, the LL estimated by the ratio copula, samples from the IGC, and the LL estimated from the TLL.
C.3 High dimensional experiments
The two high-dimensional datasets used are the Digits dataset [Alimoglu and Alpaydin, 1997] and MNIST [LeCun, 1998]. They both show handwritten digits in $8\cdot 8$ or $28\cdot 28$ format. As both these datasets represent complex dependencies, as measured using Lemma 3.2 from the main text, we use a GG+ratio copula model. We show results for a simpler (GG+Ratio) and larger version (GG+Ratio large see Tables 6 and 8), to showcase the potential for scalability of ratio copulas. In Figure 6, we show random samples from the simpler ratio copula, the simplified vine copula, the IGC and the Gaussian copula as well as true test data. We show samples from the same copula classes fitted on the MNIST data in Figure 7. Samples shown are obtained from the copula model and then transformed back to the data space using the empirical CDF as described in the preprocessing paragraph. We note that the samples shown are random across tiles, meaning that we are not showing how well a model imitates a given target digit. We are exploring the learned distribution implied by the model after training on the data samples, doing so by showing a collection of uncurated random samples from any given model.